OpenStudy (anonymous):

integrate sqrt(10^(3x))dx

OpenStudy (anonymous):

please help me solve this question

OpenStudy (anonymous):

Perhaps substitution after rewriting as \[\int\limits 10^{3x/2}dx\] let u = 3x/2, and integrate using exponential rule.

OpenStudy (anonymous):

\[I=\int\limits \sqrt{10^{3X}}dx=\int\limits10^{\frac{ 3x }{2 }}dx\] \[put ~\frac{ 3x }{2 }=t,\frac{ 3 }{ 2 }dx=dt,dx=\frac{ 2 }{ 3 }dt\] \[\int\limits a ^{x}dx=\frac{ a ^{x} }{\log a}\]

OpenStudy (anonymous):

Thank you so much both of you for helping me i really appreciate it

OpenStudy (anonymous):

yw